# Resonant spectral cascade in pulsatile flow triggered by arterial geometry

### v1

@@ Results
The simulation was performed for a representative case with a Womersley number of $\mathrm{Wo}=10$. The initial conditions and spatially-modulated coefficients for this run are presented in Figure 1. The top panel shows the initial waveform, $a_{1}(\xi,0)$, a composite of three harmonics resulting in an asymmetric profile with a dominant peak and a broader, shallower trough. The bottom panel displays the periodic functions for the dispersion coefficient, $\beta(\xi)$, and the damping coefficient, $\eta(\xi)$. Both coefficients vary in-phase with one another over two full geometric periods within the computational domain, establishing the spatially-dependent forcing of the system.
Figure 1: Initial setup for the simulation with $\mathrm{Wo}=10$. Top: The multi-harmonic initial waveform $a_{1}(\xi,0)$. Bottom: The spatially-periodic dispersion coefficient $\beta(\xi)$ (left axis, red) and damping coefficient $\eta(\xi)$ (right axis, blue).
@@
Figure 2: Primary instability diagnostics for $\mathrm{Wo}=10$. Top: Total wave energy $I_{2}(\tau)$ on a semi-logarithmic scale. Middle: Instantaneous energy growth rate $G(\tau)$. Bottom: Spectral broadening ratio $R(\tau)$, with the broadband threshold indicated by the dashed line.
@@
Figure 3: Modal energy evolution for the first five harmonics ($k_{0}$ to $5k_{0}$) for the $\mathrm{Wo}=10$ case. The energy in each mode is plotted over time on a semi-logarithmic scale.
The physical effect of the transient energy cascade on the waveform’s shape is visualized in the snapshots of Figure 4. The initially smooth profile at $\tau=0.0$ undergoes a dramatic transformation into the highly irregular waveform seen at $\tau=50.0$, which is characterized by the emergence of numerous high-frequency oscillations and sharp gradients. The subsequent snapshots at $\tau=100.0$, $150.0$, and $200.0$ illustrate a process of gradual damping, where both the overall amplitude and the fine-scale structures are progressively attenuated.
Figure 4: Waveform shape evolution at selected time snapshots for $\mathrm{Wo}=10$. The plot illustrates the development of high-frequency structures and the subsequent decay of the wave amplitude.
@@
Figure 5: Evolution of the classical KdV invariants for $\mathrm{Wo}=10$. Top: Momentum invariant $I_{1}$. Bottom: Higher-order invariant $I_{3}$. The non-constant behavior indicates a non-conservative system.
@@
Figure 6: Spatiotemporal evolution of the wave and its spectrum for $\mathrm{Wo}=10$. Top: Waveform amplitude $a_{1}(\xi,\tau)$ as a function of space and time. Bottom: Power spectrum $|\hat{a}_{1}(k,\tau)|^{2}$ as a function of wavenumber and time.
@@
Figure 7: Summary of instability metrics from the parameter sweep versus Womersley number ($\mathrm{Wo}$). The maximum late-stage growth rate ($G$, left axis, green dashed line) remains negative for all cases. The maximum transient spectral broadening ratio ($R$, right axis, purple solid line) shows a peak at an intermediate Womersley number. Shaded regions denote the regimes of exponential growth ($G>0$) and broadband spectral cascade ($R>1.5$).
@@
Figure 8: Parametric comparison of final states from the sweep. Top: Overlay of the final waveform shapes, $a_{1}(\xi,\tau_{final})$, for each Womersley number. Bottom: Corresponding overlay of the final power spectra, $|\hat{a}_{1}(k,\tau_{final})|^{2}$, on a semi-logarithmic scale.
Figure 9: Parametric structure of final states represented as heatmaps. The vertical axis in both plots is the Womersley number. Top: The final waveform amplitude as a function of spatial position $\xi$. Bottom: The final power spectrum as a function of wavenumber $k$.
@@ Discussion
First, what is the fundamental role of geometry within this theoretical framework? The system is modeled by the variable-coefficient fKdV equation, where the influence of geometry is explicitly encoded in the spatially-varying coefficients $\beta(\xi)$ and $\eta(\xi)$. A key finding, shown in Figure 5, is the non-conservation of the classical KdV invariants, $I_{1}$ and $I_{3}$. In a constant-coefficient system, these quantities would be strictly conserved, and their fluctuating behavior here is the definitive mathematical proof that the system is being actively forced by the geometry. This validates the central premise of the model: geometry is not a passive boundary condition but an active agent that continuously modulates the integral properties and energy distribution of the flow, a concept motivated by fractional modeling theory [18, 19].
@@
Given that the system is actively forced but remains globally stable—with a negative energy growth rate $G$ across all tested Womersley numbers (Figure 7)—what is the primary physical consequence of this forcing? The results demonstrate that the forcing drives a highly efficient, transient spectral cascade. This is observed directly as the rapid peak in the spectral broadening ratio $R$ (Figure 2) and is visualized as the breakdown of the initially smooth waveform (Figure 1) into a field of complex, high-frequency structures (Figure 4 and 6). This process, whereby energy is rapidly redistributed from the initial low-wavenumber harmonics to a broad continuum of modes (Figure 3), is governed by the interaction between the nonlinear steepening term, $\alpha_{0}a_{1}\partial_{\xi}a_{1}$, and the parametrically-forced dispersive term, $\beta_{0}\varepsilon_{\beta}\cos(q\xi)\partial^{3}_{\xi}a_{1}$. Unlike an integrable system which balances these effects to form stable solitary waves, the spatial variation of $\beta(\xi)$ constantly disrupts this balance, preventing the formation of stable structures and instead promoting a chaotic-like state of high complexity. This provides a clear, reduced-order physical mechanism for the geometry-induced unsteadiness reported in other computational models [3, 12], and is consistent with the non-Kolmogorov spectral cascades and near-wall coherent structures observed in our prior high-fidelity simulations of patient-specific neurovascular geometries [21, 6].
@@
It is important to acknowledge the limitations of the present model. By design, our framework isolates geometry as the sole parametric forcing agent and assumes a rigid wall. Our prior experimental work has shown that wall compliance can have a significant, often attenuating, effect on the kinetic energy cascade and near-wall velocity oscillations in aneurysm models [5, 4]. Therefore, a complete picture of hemodynamic stability requires understanding the interplay between the geometry-driven spectral cascade identified here and the dissipative or modulating effects of wall elasticity. Future work should aim to incorporate wall compliance into this reduced-order framework, potentially as a time-dependent or nonlocal damping term, to explore this competitive dynamic.
@@ Conclusion
Arterial geometry, when modeled as parametric forcing, acts as a resonant, frequency-dependent spectral catalyst. For the studied parameters, it does not cause runaway exponential instability but instead fundamentally alters the character of the flow by efficiently generating small-scale complexity, a process whose efficiency is maximized at a specific resonant Womersley number. The Womersley number plays a dual role: it tunes the system towards this spectral resonance by controlling dispersion, while simultaneously governing the ultimate dissipation of the resulting small-scale structures. While this 1D model is a simplification of the full 3D problem, it successfully isolates this fundamental physical mechanism, providing a quantitative framework for understanding how geometry can be an active and crucial participant in determining the stability and structural complexity of hemodynamic and other complex internal flows.

### v2

@@ Results
The simulation was performed for a representative case with a Womersley number of $\mathrm{Wo}=10$. The initial conditions and spatially-modulated coefficients for this run are presented in Figure 1. The top panel shows the initial waveform, $a_{1}(\xi,0)$, a composite of three harmonics resulting in an asymmetric profile with a dominant peak and a broader, shallower trough. The bottom panel displays the periodic functions for the dispersion coefficient, $\beta(\xi)$, and the damping coefficient, $\eta(\xi)$. BBoth coefficients vary in-phase with one another over two full geometric periods within the computational domain, illustrating the prescribed geometric modulation at the model level. In the numerical time integration, the spatial means $\bar{\beta}$ and $\bar{\eta}$ define the effective dispersive and dissipative parameters used in the linear propagator, so the plotted profiles represent the imposed geometric parameterization rather than a pointwise-varying operator during time stepping.
Figure 1: Initial configuration for the representative case $\mathrm{Wo}=10$. Top: Multi-harmonic initial condition $a_{1}(\xi,0)$ composed of three Fourier modes. Bottom: Prescribed spatial profiles of the dispersion coefficient $\beta(\xi)$ (red) and damping coefficient $\eta(\xi)$ (blue), illustrating the geometric modulation introduced at the model level. The numerical evolution employs the spatial means $\bar{\beta}$ and $\bar{\eta}$ in the linear operator.
@@
Figure 2: Primary instability diagnostics for $\mathrm{Wo}=10$. Top: Total wave energy $I_{2}(\tau)$ (semi-log scale), showing monotonic decay. Middle: Instantaneous logarithmic growth rate $G(\tau)$, which remains negative for all $\tau$, confirming global stability. Bottom: Spectral broadening ratio $R(\tau)$, measuring redistribution of energy toward higher wavenumbers. A sharp transient peak precedes long-term damping, indicating efficient but non-exponential spectral transfer.
@@
Figure 3: Temporal evolution of modal energies for the first five harmonics ($k_{0}$ through $5k_{0}$) at $\mathrm{Wo}=10$. Rapid initial redistribution leads to near equipartition among low-order modes, followed by coupled decay.
The physical effect of the transient energy cascade on the waveform’s shape is visualized in the snapshots of Figure 4. The initially smooth profile at $\tau=0.0$ undergoes a marked increase in short-wavelength oscillatory content seen at $\tau=50.0$, which is characterized by the emergence of numerous high-frequency oscillations and sharp gradients. The subsequent snapshots at $\tau=100.0$, $150.0$, and $200.0$ illustrate a process of gradual damping, where both the overall amplitude and the fine-scale structures are progressively attenuated.
Figure 4: Waveform evolution at selected times for $\mathrm{Wo}=10$. The initially smooth multi-harmonic profile develops short-wavelength oscillations during the transient redistribution phase, followed by gradual attenuation of both amplitude and fine-scale structure as dissipation dominates.
@@
Figure 5: Evolution of classical KdV invariants for $\mathrm{Wo}=10$. Top: Momentum-like invariant $I_{1}$. Bottom: Higher-order invariant $I_{3}$. Deviation from constant values confirms that the system is non-conservative due to dispersive and dissipative effects, in contrast to the integrable constant-coefficient KdV limit.
@@
Figure 6: Spatiotemporal evolution for $\mathrm{Wo}=10$. Top: Field $a_{1}(\xi,\tau)$ showing breakdown of large-scale coherence into small-scale structures during the transient phase. Bottom: Corresponding power spectrum $|\hat{a}_{1}(k,\tau)|^{2}$, demonstrating rapid broadening across wavenumbers before global decay reduces spectral intensity.
@@
Figure 7: Parameter sweep across $\mathrm{Wo}\in\{2,5,10,15,20\}$. Left axis: Maximum late-time growth rate $G$, which remains negative ($G<0$) for all cases, indicating absence of exponential instability. Right axis: Maximum transient spectral broadening ratio $R$, showing a non-monotonic dependence on $\mathrm{Wo}$ with a peak near $\mathrm{Wo}=15$. Shaded regions indicate regimes of exponential growth ($G>0$, not observed) and broadband redistribution ($R>1.5$).
@@
Figure 8: Final-state comparison across Womersley numbers. Top: Final waveform $a_{1}(\xi,\tau_{\mathrm{final}})$ for each $\mathrm{Wo}$. Amplitude decreases at large $\mathrm{Wo}$ as damping strengthens. Bottom: Corresponding final power spectra $|\hat{a}_{1}(k,\tau_{\mathrm{final}})|^{2}$. Intermediate $\mathrm{Wo}$ values produce broader spectra, whereas low $\mathrm{Wo}$ cases retain energy concentrated at low wavenumbers.
Figure 9: Color-coded representation of final states across Womersley numbers. Top: Final waveform amplitude as a function of spatial coordinate $\xi$ and $\mathrm{Wo}$. Bottom: Final power spectrum magnitude $|\hat{a}_{1}(k,\tau_{\mathrm{final}})|^{2}$. Spectral width increases from $\mathrm{Wo}=2$ to approximately $\mathrm{Wo}=15$, then contracts at $\mathrm{Wo}=20$ due to stronger dissipation, consistent with the non-monotonic behavior of the spectral broadening ratio.
@@ Discussion
First, what is the fundamental role of geometry within this theoretical framework? The system is modeled by the variable-coefficient fKdV equation, where the influence of geometry is explicitly encoded in the spatially-varying coefficients $\beta(\xi)$ and $\eta(\xi)$. A key finding, shown in Figure 5, is the non-conservation of the classical KdV invariants, $I_{1}$ and $I_{3}$. In the constant-coefficient KdV limit these quantities are strictly conserved, so their deviation here reflects the presence of dispersive and dissipative terms arising from geometry-parameterized coefficients. Although the numerical evolution employs the spatial means of $\beta(\xi)$ and $\eta(\xi)$ in the linear propagator, the resulting dynamics remain distinct from the integrable limit and therefore quantify the influence of geometry-induced parameter renormalization on spectral redistribution [23, 24].
Observability and experimental detectability
The geometry-induced spectral redistribution discussed here is an axial-wavenumber effect: energy is transferred from the initially imposed low-wavenumber content (set by $k_{0}$) toward higher axial wavenumbers, as quantified by the spectral broadening ratio $R(\tau)$ defined from the Fourier spectrum of $a_{1}(\xi,\tau)$. In the present nondimensionalization $\xi=x/L_{0}$, so an axial wavenumber $k$ corresponds to a dimensional wavelength $\lambda=2\pi L_{0}/k$. Accordingly, the smallest axial length scale explicitly interrogated by the diagnostic (i.e., the “high-kk” band above the cutoff $k_{c}$ used in defining $R$) is
$\lambda_{\min}\sim\frac{2\pi L_{0}}{k_{c}}.$
With the representative choices used herein ($k_{c}=O(1$$10)$), $\lambda_{\min}$ is $O(L_{0})$; for large conduit arteries a natural choice is $L_{0}\sim R_{0}$, implying $\lambda_{\min}=O(R_{0})$, i.e. millimetric-to-centimetric axial scales for $R_{0}\sim 5$1515 mm.
This scale separation indicates that the predicted redistribution is, in principle, accessible to current velocity-measurement modalities in large arteries. In vivo, 4D flow MRI provides time-resolved volumetric velocity fields and is suited to resolving axial variations at millimetric spatial resolution over the cardiac cycle; the model prediction would manifest as a transient increase in higher-$k$ content of the axial velocity (or derived wall shear stress) along the vessel centerline or along a centerline-aligned curvilinear coordinate. High-frame-rate Doppler ultrasound, while spatially more limited, directly reports spectral broadening of the velocity signal and can therefore probe time-localized changes in spectral content over the cardiac period. Finally, in vitro particle image velocimetry (PIV) in compliant or rigid phantoms can provide the spatial and temporal resolution needed to resolve the full axial spectrum and to validate the predicted redistribution mechanism under controlled geometric modulation. The present results therefore support a measurable signature at conduit-artery scales, while acknowledging that resolving substantially smaller scales than $\lambda_{\min}$ would require dedicated experimental designs beyond standard clinical protocols.
@@
Given that the system is actively forced but remains globally stable—with a negative energy growth rate $G$ across all tested Womersley numbers (Figure 7)—what is the primary physical consequence of this forcing? The results demonstrate that the forcing drives a highly efficient, transient spectral cascade. This is observed directly as the rapid peak in the spectral broadening ratio $R$ (Figure 2) and is visualized as the breakdown of the initially smooth waveform (Figure 1) into a field of complex, high-frequency structures (Figure 4 and 6). This process, whereby energy is rapidly redistributed from the initial low-wavenumber harmonics to a broad continuum of modes (Figure 3), is governed by the interaction between the nonlinear steepening term, $\alpha_{0}a_{1}\partial_{\xi}a_{1}$, and the parametrically-forced dispersive term, $\beta_{0}\varepsilon_{\beta}\cos(q\xi)\partial^{3}_{\xi}a_{1}$. Unlike an integrable system which balances these effects to form stable solitary waves, the geometry-parameterized dispersive response disrupts this balance, preventing the formation of stable structures and instead promoting a chaotic-like state of high complexity. This provides a clear reduced-order mechanism for one class of geometry-sensitive unsteady behavior reported in other computational models [3, 17], and is consistent with the non-Kolmogorov spectral cascades and near-wall coherent structures observed in our prior high-fidelity simulations of patient-specific neurovascular geometries [7, 6].
This interpretation also aligns, at a broader phenomenological level, with prior CFD studies showing that vascular geometry can reorganize unsteady pulsatile flow in ways that are strongly phase- and morphology-dependent. In idealized curved or bifurcating configurations, geometric variation has been shown to modify flow partitioning, alter oscillatory wall-shear patterns, shift stagnation dynamics, and break planar symmetry over the cardiac cycle [26, 27, 28, 29]. In more irregular or diseased geometries, numerical studies have further reported enhanced disturbance levels, transitional behavior, broadband velocity fluctuations, and Doppler-observable spectral broadening under pulsatile conditions [30, 31, 32, 33, 34, 35]. These results support the general premise that arterial geometry influences the temporal organization and frequency content of pulsatile flow. Nevertheless, those studies do not isolate the specific mechanism identified here. The role of the present model is therefore explanatory. This study extracts and dissects an interpretable mechanism by which geometry can act as a spectral modulator in the main human arteries.
Harmonic Regulation
@@
It is important to acknowledge the limitations of the present model. By design, our framework isolates geometry as the sole parametric forcing agent and assumes a rigid wall.
[…]
For large conduit arteries (e.g., aorta, renal, carotid), physiological parameters satisfy $U\ll c_{w}$, implying $\chi\ll 1$. In this strongly sub-wave-speed regime, wall motion modifies the axial momentum balance primarily through $O(\chi)$ corrections to phase speed and through additional attenuation associated with pressure–area energy exchange. Importantly, the leading-order inertial–viscous balance that generates Womersley dispersion remains dominant. Consequently, compliance acts predominantly as a renormalization of the effective dispersive and dissipative response, shifting quantitative peak timing and amplitude, rather than eliminating the geometry-modulated phase-redistribution mechanism identified here. The author’s prior experimental work has shown that wall compliance can have a significant, often attenuating, effect on the kinetic energy cascade and near-wall velocity oscillations in aneurysm models [5, 4]. Therefore, a complete picture of hemodynamic stability requires understanding the interplay between the geometry-driven spectral cascade identified here and the dissipative or modulating effects of wall elasticity. Future work should aim to incorporate wall compliance into this reduced-order framework, potentially as a time-dependent or nonlocal damping term, to explore this competitive dynamic.
@@ Conclusion
When geometry is parameterized through dispersive and dissipative coefficients within a reduced-order pulsatile-flow model, it produces a resonant, frequency-dependent redistribution of spectral energy. For the studied parameter range, this mechanism does not generate exponential instability ($G<0$ for all cases), but it significantly modifies the distribution of energy across axial wavenumbers. The efficiency of this redistribution exhibits a non-monotonic dependence on the Womersley number, with a distinct peak at intermediate values.
The Womersley number therefore plays a dual regulatory role: it governs both the intrinsic dispersive scaling and the strength of dissipation, determining whether energy remains concentrated at low wavenumber or is transiently redistributed toward smaller axial scales. While the present one-dimensional formulation does not resolve full three-dimensional hemodynamics or fluid–structure interaction, it isolates a fundamental mechanism by which geometry-dependent parameter variations can reshape spectral structure without inducing runaway instability.
These results provide a quantitative framework for interpreting how geometric complexity may influence spectral organization in pulsatile internal flows and suggest experimentally accessible diagnostics based on axial spectral redistribution.
@@
Supplementary Information
Detailed Derivation of the Fractional KdV Model
This appendix provides a graduate-level derivation of the one-dimensional variable-coefficient fractional Korteweg–de Vries (fKdV) equation used in the main text. The derivation proceeds from the incompressible Navier–Stokes equations in cylindrical coordinates, through a weakly nonlinear perturbation analysis, radial-mode projection, and non-dimensionalization, yielding a reduced-order model that incorporates geometric dispersion, nonlinear steepening, and fractional geometric damping.
A.1. Governing equations and geometric assumptions
Blood is modeled as an incompressible Newtonian fluid with density $\rho$ and dynamic viscosity $\mu$. In cylindrical coordinates $(r,\theta,z)$, the velocity field is $\mathbf{v}=(v_{r},v_{\theta},v_{z})$ and the pressure is $p$. The continuity and axial momentum equations are
$\frac{1}{r}\frac{\partial}{\partial r}(rv_{r})+\frac{1}{r}\frac{\partial v_{\theta}}{\partial\theta}+\frac{\partial v_{z}}{\partial z}$ $=0,$ (A.1)
$\rho\!\left(\frac{\partial v_{z}}{\partial t}+v_{r}\frac{\partial v_{z}}{\partial r}+\frac{v_{\theta}}{r}\frac{\partial v_{z}}{\partial\theta}+v_{z}\frac{\partial v_{z}}{\partial z}\right)$ $=-\frac{\partial p}{\partial z}+\mu\!\left[\frac{1}{r}\frac{\partial}{\partial r}\!\left(r\frac{\partial v_{z}}{\partial r}\right)+\frac{1}{r^{2}}\frac{\partial^{2}v_{z}}{\partial\theta^{2}}+\frac{\partial^{2}v_{z}}{\partial z^{2}}\right].$ (A.2)
We adopt the following assumptions, standard in pulsatile hemodynamics and long-wave asymptotics:
  • Axisymmetry: $\partial/\partial\theta=0$ and $v_{\theta}=0$;
  • Slenderness: $R_{0}/L_{0}=\delta\ll 1$ for characteristic radius $R_{0}$ and axial scale $L_{0}$;
  • Mild curvature/torsion: curvature $\kappa(x)$ and torsion $\tau(x)$ satisfy $\kappa R_{0},\tau R_{0}\ll 1$ and act through slow axial modulation of coefficients.
Under axisymmetry, (A.2) simplifies to
$\rho\!\left(\frac{\partial v_{z}}{\partial t}+v_{r}\frac{\partial v_{z}}{\partial r}+v_{z}\frac{\partial v_{z}}{\partial z}\right)=-\frac{\partial p}{\partial z}+\mu\!\left[\frac{1}{r}\frac{\partial}{\partial r}\!\left(r\frac{\partial v_{z}}{\partial r}\right)+\frac{\partial^{2}v_{z}}{\partial z^{2}}\right].$ (A.3)
A.2. Base oscillatory Womersley flow
The zeroth-order solution is the classical Womersley flow driven by a harmonic pressure gradient
$\frac{\partial p_{0}}{\partial z}=\Re\left\{\hat{P}(x)e^{i\omega t}\right\}.$ (A.4)
Under Womersley conditions, the convective terms vanish identically:
$\rho\frac{\partial v_{0}}{\partial t}=-\frac{\partial p_{0}}{\partial z}+\mu\left[\frac{1}{r}\frac{\partial}{\partial r}\!\left(r\frac{\partial v_{0}}{\partial r}\right)\right].$ (A.5)
Its solution is the known Womersley profile
$v_{0}(r,t;x)=\Re\!\left[\frac{\hat{P}(x)}{i\rho\omega}\left(1-\frac{J_{0}(i^{3/2}\mathrm{Wo}\,r/R_{0})}{J_{0}(i^{3/2}\mathrm{Wo})}\right)e^{i\omega t}\right],$ (A.6)
where $\mathrm{Wo}=R_{0}\sqrt{\omega/\nu}$ is the Womersley number, $\nu=\mu/\rho$.
We seek an evolution equation for perturbations $v^{\prime}$, writing
$v_{z}(r,x,t)=v_{0}(r,t;x)+v^{\prime}(r,x,t).$ (A.7)
A.3. Perturbation ansatz and long-wave scaling
We introduce a small parameter $\varepsilon\ll 1$ and expand the perturbation in the leading radial eigenmode $\phi(r)$:
$v^{\prime}(r,x,t)=\varepsilon A(x,t)\phi(r)+O(\varepsilon^{2}).$ (A.8)
We assume Womersley oscillations act on a fast timescale $t_{0}=t$ while the perturbation evolves on a slow timescale $t_{1}=\varepsilon t$. Using long-wave scaling,
$\frac{\partial}{\partial x}\sim O(\varepsilon),\quad\frac{\partial}{\partial t_{1}}\sim O(\varepsilon),$
we let derivatives transform as
$\partial_{t}\to\partial_{t_{0}}+\varepsilon\partial_{t_{1}}.$ (A.9)
A.4. Radial eigenproblem and projection
Expanding the pressure as $p=p_{0}+\varepsilon p_{1}+\cdots$, the $O(\varepsilon)$ perturbation satisfies
$\rho\phi\,\frac{\partial A}{\partial t_{0}}=-\frac{\partial p_{1}}{\partial z}+\mu A\mathcal{L}_{r}[\phi],\qquad\mathcal{L}_{r}[\phi]=\frac{1}{r}\frac{d}{dr}\!\left(r\frac{d\phi}{dr}\right).$ (A.10)
We introduce the cross-sectional inner product
$\langle f,g\rangle=\int_{0}^{R_{0}}f(r)g(r)\,r\,dr,$ (A.11)
and normalize $\phi$ such that $\langle\phi,\phi\rangle=1$.
Projecting the full $O(\varepsilon^{2})$ equation onto $\phi(r)$ produces a solvability condition yielding a scalar amplitude equation of the form
$\frac{\partial A}{\partial t}+C(x)A\frac{\partial A}{\partial x}+D(x)\frac{\partial^{3}A}{\partial x^{3}}+K(x)\mathcal{D}^{\gamma(x)}A=0,$ (A.12)
where $C(x)$, $D(x)$, and $K(x)$ are determined by radial integrals involving $\phi$ and geometric corrections.
A.5. Nonlinear term
The leading nonlinearity arises from
$v^{\prime}\frac{\partial v^{\prime}}{\partial z}=\varepsilon^{2}A\,\frac{\partial A}{\partial x}\phi^{2}(r)+\cdots$
and additional contributions involving $v_{r}\partial_{r}v^{\prime}$. After projection, this yields a term
$C(x)A\frac{\partial A}{\partial x},$ (A.13)
with
$C(x)\sim\rho\int_{0}^{R_{0}}\!\phi^{2}(r)\phi(r)\,r\,dr,$
up to geometry-induced corrections.
A.6. Dispersive term
Long-wave axial variations and geometry introduce higher-order axial derivatives. The leading dispersive correction takes the form
$D(x)\frac{\partial^{3}A}{\partial x^{3}},$ (A.14)
where $D(x)$ depends on viscosity, curvature, torsion, and Womersley-dependent radial coupling.
A.7. Fractional geometric damping
Geometric roughness, tortuosity, and distributed compliance introduce multiscale nonlocal damping. This is modeled by a Riesz fractional operator
$\mathcal{F}\{\mathcal{D}^{\gamma}A\}(k)=|k|^{1+\gamma}\hat{A}(k),$ (A.15)
giving the term
$K(x)\mathcal{D}^{\gamma(x)}A.$ (A.16)
In this work, we set $\gamma(x)=\gamma_{0}=0$ to isolate geometric forcing through $D(x)$ and $K(x)$.
A.8. Non-dimensionalization
Introduce
$\xi=\frac{x}{L_{0}},\qquad\tau=\omega t,\qquad A(x,t)=A_{0}a_{1}(\xi,\tau),$
so that
$\partial_{t}=\omega\partial_{\tau},\quad\partial_{x}=\frac{1}{L_{0}}\partial_{\xi},\quad\partial_{x}^{3}=\frac{1}{L_{0}^{3}}\partial_{\xi}^{3}.$
Equation (A.12) becomes
$\frac{\partial a_{1}}{\partial\tau}+\alpha(\xi)a_{1}\frac{\partial a_{1}}{\partial\xi}+\beta(\xi)\frac{\partial^{3}a_{1}}{\partial\xi^{3}}+\eta(\xi)(-\partial_{\xi}^{2})^{\frac{1+\gamma_{0}}{2}}a_{1}=0,$ (A.17)
with
$\alpha(\xi)=\frac{C(x)A_{0}}{\omega L_{0}},\quad\beta(\xi)=\frac{D(x)}{\omega L_{0}^{3}},\quad\eta(\xi)=\frac{K(x)}{\omega L_{0}^{1+\gamma_{0}}}.$ (A.18)
We choose $A_{0}$ such that $\alpha(\xi)\equiv 1$.
A.9. Geometric modulation
Geometry modulates dispersion and damping through
$\beta(\xi)=\beta_{0}\big[1+\varepsilon_{\beta}\cos(q\xi)\big],\qquad\eta(\xi)=\eta_{0}\big[1+\varepsilon_{\eta}\cos(q\xi)\big],$ (A.19)
where $q$ is the geometric wavenumber and $\varepsilon_{\beta},\varepsilon_{\eta}\leq 0.3$.
A.10. Scaling of the dispersion coefficient with Womersley number
Dimensional consistency of dispersive corrections in oscillatory pipe flow suggests
$D\propto\nu^{2}\mathrm{Wo}^{2},$
while the pulsation time scale satisfies
$T_{0}=\omega^{-1}=\frac{R_{0}^{2}}{\nu\mathrm{Wo}^{2}}.$
Thus
$\beta_{0}=\frac{DT_{0}}{L_{0}^{3}}\propto\frac{\nu R_{0}^{2}}{L_{0}^{3}}\mathrm{Wo}^{-2}.$
We therefore adopt
$\beta_{0}=\beta_{\mathrm{ref}}\,\mathrm{Wo}^{-2},$ (A.20)
consistent with decreasing dispersive coupling at high Womersley number.
Damping is modeled as
$\eta_{0}=\eta_{\mathrm{ref}}\left(1+\frac{0.1}{\mathrm{Wo}}\right).$ (A.21)
The form $\eta_{0}=\eta_{\mathrm{ref}}(1+C/\mathrm{Wo})$ follows from oscillatory boundary-layer scaling. The Stokes boundary-layer thickness associated with pulsatile motion is $\delta\sim\sqrt{2\nu/\omega}$, hence $\delta/R_{0}=O(\mathrm{Wo}^{-1})$. Since near-wall shear-driven dissipation scales with the boundary-layer contribution, the leading frequency-dependent correction to an effective damping coefficient is proportional to $\delta/R_{0}$, motivating $C=O(10^{-1})$ as an order-unity prefactor smaller than one. I choose $C=0.1$ so that this correction is weak in the large-artery regime emphasized in the manuscript (i.e., it remains a small perturbation for $\mathrm{Wo}\gtrsim 5$) while preserving the correct asymptotic trend as $\mathrm{Wo}$ decreases. The constant $0.1$ is therefore a conservative phenomenological coefficient rather than a claimed universal constant.
A.11. Limiting cases
  • KdV limit: $\varepsilon_{\beta}=\varepsilon_{\eta}=0$ and $\eta_{0}=0$ gives
    $a_{1,\tau}+a_{1}a_{1,\xi}+\beta_{0}a_{1,\xi\xi\xi}=0.$
  • Fractional Burgers limit: $\beta_{0}=0$ yields
    $a_{1,\tau}+a_{1}a_{1,\xi}+\eta(\xi)(-\partial_{\xi}^{2})^{(1+\gamma_{0})/2}a_{1}=0.$
  • No geometry limit: $\varepsilon_{\beta}=\varepsilon_{\eta}=0$ reduces the system to a constant-coefficient equation in which spatial parametric forcing is absent. Frequency-dependent optimal spectral redistribution may still occur through the Womersley scaling of $\beta_{0}$ and $\eta_{0}$.
A.12. Final reduced model
The final dimensionless model is
$\boxed{\frac{\partial a_{1}}{\partial\tau}+a_{1}\frac{\partial a_{1}}{\partial\xi}+\beta(\xi)\frac{\partial^{3}a_{1}}{\partial\xi^{3}}+\eta(\xi)(-\partial_{\xi}^{2})^{1/2}a_{1}=0}$ (A.22)
with $\beta(\xi)$ and $\eta(\xi)$ given by (A.19). This is the variable-coefficient fractional KdV model governing the amplitude of weakly nonlinear perturbations to pulsatile Womersley flow in geometrically complex arteries. In the weak-modulation, zero-mean case considered in the main manuscript, this variable-coefficient model admits a leading-order homogenized effective-medium limit, described next, which is the form evolved by the numerical solver.
A.13. Homogenized effective-medium limit and numerical implementation
The spatially modulated coefficients in (A.22) may be decomposed as
$\beta(\xi)=\bar{\beta}+\tilde{\beta}(\xi),\qquad\eta(\xi)=\bar{\eta}+\tilde{\eta}(\xi),$
where the overbar denotes the spatial mean over one geometric period and $\langle\tilde{\beta}\rangle=\langle\tilde{\eta}\rangle=0.$
For the cosine modulation (A.19) with $\varepsilon_{\beta},\varepsilon_{\eta}\leq 0.3$, the fluctuating components satisfy $|\tilde{\beta}|/\bar{\beta}=O(\varepsilon_{\beta})$, $|\tilde{\eta}|/\bar{\eta}=O(\varepsilon_{\eta})$.
Substituting this decomposition into (A.22) gives
$a_{1,\tau}+a_{1}a_{1,\xi}+\bar{\beta}a_{1,\xi\xi\xi}+\bar{\eta}(-\partial_{\xi}^{2})^{1/2}a_{1}=-\tilde{\beta}(\xi)a_{1,\xi\xi\xi}-\tilde{\eta}(\xi)(-\partial_{\xi}^{2})^{1/2}a_{1}.$
The right-hand side represents a zero-mean parametric perturbation of order $O(\varepsilon)$. For weak modulation and long-wave scaling, these terms enter at higher order in the averaged phase-dispersion balance and o not alter the leading-order, spatially averaged dispersion relation that governs the dominant phase-dispersion balance in the weak-modulation regime considered in this work.
Accordingly, the numerical simulations in the main manuscript evolve the homogenized leading-order equation
$a_{1,\tau}+a_{1}a_{1,\xi}+\bar{\beta}a_{1,\xi\xi\xi}+\bar{\eta}(-\partial_{\xi}^{2})^{1/2}a_{1}=0,$
which corresponds to the effective-medium limit of the weakly modulated geometry model.
The reported non-monotonic Womersley dependence therefore reflects the frequency-dependent dispersive–dissipative balance of the homogenized operator, while the geometric modulation parameterization specifies the admissible variability bounds and provides the physiological interpretation of the effective coefficients.